OpenStudy (anonymous):

A body moves at a speed of 100 m/s for 10 sec and then moves at a speed of 20 m/s for 20 sec along the same direction. The average speed of the body is????

OpenStudy (anonymous):

using speed = distance / time distance= speed* time distance 1 = 100* 10= 1000m distance2 = 20* 20= 400m total distance= 1400 m total time = 30 sec so av. speed= 1400/30=46.67 m.s